GSFTOPK(1) GSFTOPK(1) NAME gsftopk−renderaghostscriptfontinTeXpkformat SYNOPSIS gsftopk[−ipath][−q][−t][−−debug=n][−−dosnames][−−interpreter=path][−−mapline=line] [−−mapfile=file][−−quiet][−−test][−−help][−−version]fontdpi ARGUMENTS fontNameofthefonttobecreated. dpiDesiredresolutionofthefonttobecreated,indotsperinch.Thismaybearealnumber. DESCRIPTION gsftopkisaprogramwhichcallsuptheghostscriptprogramgs(1)torenderagivenfontata givenresolution.Itpackstheresultingcharactersintothepkfileformatandwritesthemtoafile whosenameisformedfromthefontnameandtheresolution(roundedtothenearestinteger). ThefontmaybeinanyformatacceptabletoGhostscript,including.pfa,.pfb,.gsf,and.ttffiles. Thisprogramshouldnormallybecalledbyascript,suchasmktexpk,tocreatefontsondemand. gsftopkobtainsthecharacterwidthsfromthe.tfmfile,whichmustexistinthestandardsearch path.Italsomustbeabletofindthefontinamapfile(suchaspsfonts.map),formattedasin dvips(1)),unlessthe−−maplineoptionisused.Thesetofmapfilesisgivenbythe−−mapfileoption,orinthefilesconfig.ps,$HOME/.dvipsrc,andconfig.gsftopk(aswouldbeused bydvips-Pgsftopk). Thefollowingpk"specials"areaddedattheendoftheoutputfile,toprovideaninternalcheck onthecontentsofthefile:"jobname=font","mag=1","mode=modeless",and"pixelsperinch=dpi".ThisisinaccordancewiththeTeXDirectoryStandard(TDS). OPTIONS −−debug=n SettheKpathseadebugflagsaccordingtotheintegern. −−dosnames Useanameoftheformfont.pkinsteadoffont.dpipk. −h,−−help Printabriefhelpsynopsisandexit. −ipath,−−interpreter=path UsepathastheGhostscriptinterpreter. −−mapfile=file Usefiletolookforthemapinformationforfont.Thisshouldbethefullnameofthefile (inotherwords,nopathsearchingalgorithmsareapplied). −−mapline=line Uselineinsteadoflookingforanentryinamapfile.Thefirstwordoflinemustmatch font. −q,−−quiet Operatequietly;i.e.,withoutwritinganymessagestothestandardoutput. −t,−−test Testrun:returnzerostatusifthefontcanbefoundinthemapfile(s),andnonzerostatusifitcannot.Ifthisoptionisspecified,thenthedpiargumentisoptional(sincethe fontwillnotbegenerated). −v,−−version Printtheversionnumberandexit. ENVIRONMENTVARIABLES TFMFONTSColon-separatedlistofpathstosearchforthe.tfmfileassociatedwith thefont.Anextracoloninthelistwillincludethecompiled-indefault 28November1998 1 GSFTOPK(1) GSFTOPK(1) pathsatthatpoint.Adoubleslashwillenablerecursivesubdirectory searchingatthatpointinthepath. DVIPSHEADERSColon-separatedlistofpathstosearchfortheGhostscriptdriverfile render.psandforanyPostScriptheaderorfontfiles(.enc,.pfa,.pfb, .gsf,or.ttffiles).DoubleslashesandextracolonsbehaveaswithTFM-FONTS. TEXCONFIGColon-separatedlistofpathstosearchformapfiles.Doubleslashesand extracolonsbehaveaswithTFMFONTS. DVIPSRCNameoffiletoreadinsteadof$HOME/.dvipsrc.Thisshouldbethe fullnameofthefile(inotherwords,nopathsearchingalgorithmsare applied). CONFIGURATION InordertodeterminethesetofmapfilestobeusedandthepathforfindingPostScriptfiles, gsftopkreads,inorder,thefilesconfig.ps,.dvipsrc,andconfig.gsftopk.Thefilesconfig.ps andconfig.gsftopkaresearchedforusingtheenvironmentvariableTEXCONFIG,theKpathseaconfigurationfile,orthecompiled-indefaultpaths.Thefile.dvipsrcissearchedforinthe user’shomedirectory. Thesefilesareinthesameformatasfordvips(aswellasbeinginthesamelocations).The entriesusedbygsftopkareasfollows. HpathIndicatesthattheGhostscriptdriverfilerender.psandthePostScriptheaderandfont filesaretobesearchedforusingpath. pfileIndicatesthatthelistofmapfilesistobeerasedandreplacedbyfile. p+fileIndicatesthatfileistobeaddedtothelistofmapfiles. Allotherentriesareignored. Thisissimilartothehandlingoftheseoptionswhenrunningdvips-Pgsftopk.Formore details,seetheKpathseamanual. BUGS gsftopksometimeshastroublewithfontswithverycomplicatedcharacters(suchastheSealof theUniversityofCalifornia).Thisisbecausegsftopkusesthecharpathoperatortodetermine theboundingboxofeachcharacter.Ifthecharacteristoocomplicated,thenoldversionsof Ghostscriptfail,causinggsftopktoterminatewithanerrormessage Calltogsstoppedbysignal10 (Thenumbermayvaryfromsystemtosystem;itcorrespondstoabuserrororasegmentation fault.)Thebestwaytofixthisbugistoinstallacurrentversionofghostscript.Asanalternative,gsftopkcanbeinstructedtousetheboundingboxprovidedwiththefont(ifoneexists) insteadoffindingaboundingboxforeachcharacter.Todothis,includethestring /usefontbboxtruedef inthefontmapfile;e.g., ucseal"/usefontbboxtruedef" Thiswillnotaffectuseofthefontbydvips. SEEALSO gs(1),gftopk(1),tex(1),xdvi(1),dvips(1) AUTHOR WrittenbyPaulVojta.ThisprogramwasinspiredbyKarlBerry’sgsrenderfont. MODIFICATIONS ModifiedbyYvesArrouyetouseKarlBerry’sKpathsealibrary. 28November1998 2